STIMULATED EMISSION PUMPING SPECTROSCOPY OF THE FIRST EXCITED SINGLET STATE OF GERMYLIDENE (H_{2}C=Ge)

\underline{SHENG-GUI HE} and DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

The Ã ^{1}A_{2} states of H_{2}CGe and D_{2}CGe have been explored for the first time by Ã - X̃ laser-induced fluorescence (LIF) spectroscopy of the orbitally forbidden S_{1} - S_{0} transition and stimulated emission pumping (SEP) and wavelength resolved fluorescence studies of the allowed B̃ - Ã electronic transition. Medium-resolution SEP studies gave the excited Ã state ν_{2}, ν_{3}, ν_{4}, and ν_{6} vibrational frequencies for H_{2}C^{74}Ge and D_{2}C^{74}Ge. The 4^{1} and 6^{1} levels and higher combination and overtone states are strongly Coriolis coupled, which perturbs the rotational subband structure, limiting the accuracy of the determination of the vibrational frequencies. High-resolution SEP studies of the B̃ - Ã 0_{0}^{0} band have allowed us to determine the rotational constants of the Ã state of H_{2}C^{74}Ge, from which we were able to calculate an approximate r_{0} structure with the CH bond length constrained to the ground state value. The zero-point level of D_{2}C^{74}Ge is substantially perturbed, most plausibly by interaction with an excited vibrational level of the nearby triplet (ã ^{3}A_{2}) state.